# “How Big is Big Enough?” Adjusting Model Size in Continual Gaussian Processes

### v1

@@ F.3 UCI datasets
Figure 9: A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. We show the estimate of the magnitude field learned sequentially after travelling the path shown in a dotted line. The degree of transparency represents the marginal variance.
F.4 Magnetic anomalies
The data used in this experiment is obtained from Solin et al. [39] and can be accessed at https://github.com/AaltoML/magnetic-data. The objective of this task is to detect local anomalies in the Earth’s magnetic field online, caused by the presence of bedrock and magnetic materials in indoor building structures. For this purpose, a small robot with a 3-axis magnetometer moves around an indoor space of approximately 6 meters by 6 meters and measures the magnetic field strength. Out of the 9 available trajectories, we use trajectories 1, 2, 3, 4, and 5 (with $n=8875,9105,9404,7332,8313$, respectively) for the experiments. Specifically, we use trajectories 1, 2, 4 and 5 for the first experiment and trajectory 3 for the second.
In this study, we use the setup proposed in Chang et al. [4]. The proposed model applies a GP prior to magnetic field strength, given by $\mathcal{G}\mathcal{P}\left(0,\sigma_{0}^{2}+\kappa_{\sigma^{2},\ell}^{\mathrm{Mat}}\left(\mathbf{x},\mathbf{x}^{\prime}\right)\right)$ (in $\mu\mathrm{T}$ ), where the kernel consist of a constant kernel and a Matérn-$\nu/2$ kernel. The model assumes the spatial domain is affected by Gaussian noise with a variance $\sigma_{\mathrm{n}}^{2}$. The initial variance for the constant kernel is set to 500, and the Gaussian likelihood is initialised with a noise variance of 0.1. Different to their setup, we do not restrict the number of inducing points, as in the rest of the experiments we set $\delta=0.05$ for VIPS.
In the first experiment, we aim to sequentially learn the paths taken by the robot using trajectories 1, 2, 4 and 5, i.e. an entire path will correspond to one batch. We investigate whether our method can adapt to changes in the environment and adjust the number of inducing points accordingly. During this process, we concurrently learn the hyperparameters $\sigma_{0}^{2}$, $\sigma^{2}$, $\ell$, and $\sigma_{\mathrm{n}}^{2}$. Our method enables us to reconstruct a representation of the magnetic field without the need to specify the number of inducing points. Figure 9 shows the temporally updating field estimate over batches alongside the corresponding path travelled in each batch.
In the second experiment, we focus on the streaming learning of trajectory 3. The trajectory is split into 20 batches. We compare our method VIPS to Conditional Variance. Since the setting in this experiment mimics an ever-expanding domain, different from Chang et al. [4] we do not restrict the number of inducing points for Conditional Variance and set $\eta=2.2$ for the whole task. This tolerance was chosen for Conditional Variance to achieve the same performance as VIPS with a prudent number of inducing points. Conditional Variance can expand its capacity as it encounters more data, however setting a fixed value for $\eta$ is a challenge, since this will depend on dataset properties. In contrast, a single parameter setting works well for our method VIPS. Note that we were able to reconstruct a similar local estimate of the first experiment, showing the versatility of our methods to different continual learning scenarios and data distributions. The comparison with OIPS is omitted since the method was not able to add sufficient inducing points to learn the path. Detailed learning of the path for each method is shown in figures 7 and 8.
Appendix G Derivation the optimal form of variational distribution
These derivations follow Bui et al. [2]. We start by rewriting the online lower bound,
@@ Appendix G Derivation the optimal form of variational distribution
$\widehat{\mathcal{L}}$ $=-\int\text{d}fq_{new}(f)\left[\log\frac{p(\mathbf{a}|\theta_{old})q_{new}(\mathbf{b})}{p(\mathbf{b}|\theta_{new})q_{old}(\mathbf{a})p(\mathbf{y}_{new}|f)}\right]$ (28)
$=\int\text{d}fq_{new}(f)\log p(\mathbf{y}_{new}|f)-\int\text{d}fq_{new}(f)\log\frac{p(\mathbf{a}|\theta_{old})q_{new}(\mathbf{b})}{p(\mathbf{b}|\theta_{new})q_{old}(\mathbf{a})}.$
@@
We can find the optimal $q_{new}$ by setting the derivative of $\widehat{\mathcal{L}}$ with respect to $q_{new}(\mathbf{b})$ to zero,
@@
$\frac{\mathrm{d}\widehat{\mathcal{L}}}{\mathrm{d}q(\mathbf{b})}=\int\mathrm{d}fp\left(f_{\neq\mathbf{b}}\mid\mathbf{b}\right)\left[\log p(\mathbf{y}_{new}\mid\mathbf{f})-\log\frac{p\left(\mathbf{a}\mid\theta_{\text{old }}\right)q_{new}(\mathbf{b})}{p\left(\mathbf{b}\mid\theta_{\text{new }}\right)q_{old}(\mathbf{a})}\right]-1=0,$ (29)
@@
$q_{\text{opt }}(\mathbf{b})\propto p(\mathbf{b}|\theta_{new})\exp\left(\int\mathrm{d}\mathbf{a}p(\mathbf{a}\mid\mathbf{b},\theta_{new})\log\frac{q_{old}(\mathbf{a})}{p\left(\mathbf{a}\mid\theta_{\text{old }}\right)}+\int\mathrm{d}\mathbf{f}p(\mathbf{f}\mid\mathbf{b},\theta_{new})\log p(\mathbf{y}_{new}\mid\mathbf{f})\right).$
@@
$\mathrm{E}_{1}$ $=\int\operatorname{d}\mathbf{a}p(\mathbf{a}\mid\mathbf{b},\theta_{new})\log\frac{q(\mathbf{a})}{p\left(\mathbf{a}\mid\theta_{\mathrm{old}}\right)}$ (30)
@@
$=\frac{1}{2}\int\mathrm{d}\mathbf{a}\mathcal{N}\left(\mathbf{a};\mathbf{K}_{\mathbf{ab}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\mathbf{Q}_{\mathbf{a}}\right)\left(-\log\frac{\left|\mathbf{S}_{\mathbf{a}}\right|}{\left|\mathbf{K}_{\mathbf{aa}}^{\prime}\right|}-\mathbf{m}_{\mathbf{a}}^{\top}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}+\mathbf{a}^{\top}\left(\mathbf{K}_{\mathbf{aa}}^{\prime-1}-\mathbf{S}_{\mathbf{aa}}^{-1}\right)\mathbf{a}\right)$
$=\frac{1}{2}\int\mathrm{d}\mathbf{a}\mathcal{N}\left(\mathbf{a};\mathbf{K}_{\mathbf{ab}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\mathbf{Q}_{\mathbf{a}}\right)\left(-\log\frac{\left|\mathbf{S}_{\mathbf{a}}\right|}{\left|\mathbf{K}_{\mathbf{aa}}^{\prime}\right|}-\mathbf{m}_{\mathbf{a}}^{\top}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}-\mathbf{a}^{\top}\mathbf{D}_{\mathbf{a}}^{-1}\mathbf{a}\right)$
@@
$\mathrm{E}_{2}$ $=\int\mathrm{d}\mathbf{f}p(\mathbf{f}\mid\mathbf{b},\theta_{new})\log p(\mathbf{y}_{new}\mid\mathbf{f})$
@@
where $\Delta_{\mathbf{a}}$ is a constant term with respesct to $\mathbf{b}$ and $M_{\mathbf{a}}$ is the dimension of $\mathbf{a}$. The optimal form of $q_{new}$ is then given by:
$q_{\mathrm{opt}}(\mathbf{b})\propto p(\mathbf{b}|\theta_{new})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)$ (31)
@@
$q_{\mathrm{opt}}(\mathbf{b})=\frac{p(\mathbf{b}|\theta_{new})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)}{\int\mathrm{d}\mathbf{b}p(\mathbf{b}|\theta_{new})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)}$ (33)
@@
where $\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}$ is the Nyström approximation of $\mathbf{K}_{\mathbf{\hat{f}\hat{f}}}$. The trace will be small when $\mathbf{b}=\{f(\mathbf{x}_{new},\mathbf{a})\}$ and can be simplified as follows:
@@
Appendix H Online Upper Bound Implementation

### v4

@@ D.4 Magnetic anomalies
Figure 7: VIPS (Ours). A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. We show the estimate of the magnitude field learned sequentially after travelling the path shown in a dotted line. The degree of transparency represents the marginal variance.
Batch 1, M = 5.
Batch 5, M = 32.
Batch 10, M = 66.
Batch 15, M = 101.
Batch 20, M = 134.
Figure 8: VIPS (Ours). A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. Data is collected continuously as the robot moves along the path. The inducing points are represented as black dots and the line represents the travelled part of the path. We indicate the batch number and number of inducing points (M). Final RMSE = 7.55.
After path #1
After path #2
After path #4
After path #5
Figure 9: Conditional Variance. A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. We show the estimate of the magnitude field learned sequentially after travelling the path shown in a dotted line. The degree of transparency represents the marginal variance.
Batch 1, M = 12.
Batch 5, M = 597.
Batch 10, M = 2206.
Batch 15, M = 4071.
Batch 20, M = 5785.
Figure 10: Conditional Variance. A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. Data is collected continuously as the robot moves along the path. The inducing points are represented as black dots and the line represents the travelled part of the path. We indicate the batch number and number of inducing points (M). Final RMSE = 10.66.
After path #1
After path #2
After path #4
After path #5
Figure 11: OIPS. A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. We show the estimate of the magnitude field learned sequentially after travelling the path shown in a dotted line. The degree of transparency represents the marginal variance.
Batch 1, M = 10.
Batch 5, M = 27.
Batch 10, M = 43.
Batch 15, M = 59.
Batch 20, M = 76.
Figure 12: OIPS. A small robot with wheels is used to perform sequential estimation of magnetic field anomalies. Data is collected continuously as the robot moves along the path. The inducing points are represented as black dots and the line represents the travelled part of the path. We indicate the batch number and number of inducing points (M). Final RMSE = 9.42.
D.5 Model size and Data Distribution with Larger Data
We reproduce the three continual learning scenarios from App. D.1, increasing the dataset size to 50 0005000050\,00050 000 data points and the input dimensionality to three, and compare how the three methods (VIPS, OIPS and Conditional Variance) scale in this setting. This setup aims to simulate practical scenarios where an exact GP would be infeasible, and a sequential approach is required instead. We use a Squared Exponential kernel initialised with lengthscale 1.0 and variance 1.0. The noise variance is set to 0.2. We evaluate all methods at the operating points corresponding to the 10% RMSE thresholds determined from the UCI experiments (App. D.3). The data is divided into $50$ batches of $1000$ points each.
Dataset 1:
Data points are sampled uniformly from $[0,20]^{3}$, with batches ordered such that new regions of the input space are discovered gradually. Model size is expected to grow linearly as more data is observed.
Dataset 2:
Data points are again sampled uniformly from $[0,20]^{3}$, but batches are shuffled at random. This represents a setting where the entire input space is observed from the start. As a result, the novelty of each batch decreases over time, and the model size is expected to plateau.
Dataset 3:
Most data is concentrated in a narrow region, with occasional outliers. Ninety per cent of the points are drawn from a truncated Cauchy distribution on $[5,15]^{3}$ centred at 10, and the remainder uniformly from $[0,40]^{3}$. This mimics a scenario where data is dense but includes some outliers. Therefore, the model size should remain small until the region with the outliers is observed.
Figure 13 shows how the number of inducing points evolves as more data is observed for each method across the three datasets. VIPS adapts across all settings, increasing model size only when needed. OIPS shows similar trends but typically adds up to twice as many inducing points as VIPS. CV grows rapidly even when unnecessary (e.g. Dataset 2). Note that for CV we impose a cap of $5000$ inducing points.
The datasets were designed to be well-approximated with relatively few inducing points, and all methods attain near-exact accuracy. These results show that even at a larger scale, VIPS maintains a compact model, using fewer than a thousand inducing points in all cases. In contrast, the results for CV and OIPS suggest that these methods may require hyperparameter tuning to control model growth in larger settings.
Figure 13: Number of inducing points used by each method as a function of cumulative data size, across the three synthetic datasets. VIPS consistently uses fewer inducing points and only increases model size as needed. OIPS follows a similar trend but typically uses twice as many inducing points as VIPS. CV grows unnecessarily and reaches the imposed cap of 5000 points.
Appendix E Derivation the optimal form of variational distribution
The derivations follow Bui et al. (2017). We start by rewriting the online lower bound,
@@ Appendix E Derivation the optimal form of variational distribution
$\widehat{\mathcal{L}}$ $=-\int\text{d}fq_{n}(f)\left[\log\frac{p(\mathbf{a}|\theta_{o})q_{n}(\mathbf{b})}{p(\mathbf{b}|\theta_{n})q_{o}(\mathbf{a})p(\mathbf{y}_{n}|f)}\right]$ (20)
$=\int\text{d}fq_{n}(f)\log p(\mathbf{y}_{n}|f)-\int\text{d}fq_{n}(f)\log\frac{p(\mathbf{a}|\theta_{o})q_{n}(\mathbf{b})}{p(\mathbf{b}|\theta_{n})q_{o}(\mathbf{a})}.$
@@
We can find the optimal $q_{n}$ by setting the derivative of $\widehat{\mathcal{L}}$ with respect to $q_{n}(\mathbf{b})$ to zero,
@@
$\frac{\mathrm{d}\widehat{\mathcal{L}}}{\mathrm{d}q(\mathbf{b})}=\int\mathrm{d}fp\left(f_{\neq\mathbf{b}}\mid\mathbf{b}\right)\left[\log p(\mathbf{y}_{n}\mid\mathbf{f})-\log\frac{p\left(\mathbf{a}\mid\theta_{\text{old }}\right)q_{n}(\mathbf{b})}{p\left(\mathbf{b}\mid\theta_{\text{new }}\right)q_{o}(\mathbf{a})}\right]-1=0,$ (21)
@@
$q_{\text{opt }}(\mathbf{b})\propto p(\mathbf{b}|\theta_{n})\exp\left(\int\mathrm{d}\mathbf{a}p(\mathbf{a}\mid\mathbf{b},\theta_{n})\log\frac{q_{o}(\mathbf{a})}{p\left(\mathbf{a}\mid\theta_{\text{old }}\right)}+\int\mathrm{d}\mathbf{f}p(\mathbf{f}\mid\mathbf{b},\theta_{n})\log p(\mathbf{y}_{n}\mid\mathbf{f})\right).$
@@
$\mathrm{E}_{1}$ $=\int\operatorname{d}\mathbf{a}p(\mathbf{a}\mid\mathbf{b},\theta_{n})\log\frac{q(\mathbf{a})}{p\left(\mathbf{a}\mid\theta_{\mathrm{o}}\right)}$ (22)
@@
$=\frac{1}{2}\int\mathrm{d}\mathbf{a}\mathcal{N}\left(\mathbf{a};\mathbf{K}_{\mathbf{ab}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\mathbf{Q}_{\mathbf{a}}\right)\left(-\log\frac{\left|\mathbf{S}_{\mathbf{a}}\right|}{\left|\mathbf{K}_{\mathbf{aa}}^{\prime}\right|}-\mathbf{m}_{\mathbf{a}}^{\top}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}+\mathbf{a}^{\top}\left(\mathbf{K}_{\mathbf{aa}}^{\prime-1}-\mathbf{S}_{\mathbf{aa}}^{-1}\right)\mathbf{a}+2\,\mathbf{a}^{\top}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}\right)$
$=\frac{1}{2}\int\mathrm{d}\mathbf{a}\mathcal{N}\left(\mathbf{a};\mathbf{K}_{\mathbf{ab}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\mathbf{Q}_{\mathbf{a}}\right)\left(-\log\frac{\left|\mathbf{S}_{\mathbf{a}}\right|}{\left|\mathbf{K}_{\mathbf{aa}}^{\prime}\right|}-\mathbf{m}_{\mathbf{a}}^{\top}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}-\mathbf{a}^{\top}\mathbf{D}_{\mathbf{a}}^{-1}\mathbf{a}+2\,\mathbf{a}^{\top}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}\right)$
@@
$\mathrm{E}_{2}$ $=\int\mathrm{d}\mathbf{f}p(\mathbf{f}\mid\mathbf{b},\theta_{n})\log p(\mathbf{y}_{n}\mid\mathbf{f})$
@@
where $\Delta_{\mathbf{a}}$ is a constant term with respesct to $\mathbf{b}$ and $M_{\mathbf{a}}$ is the dimension of $\mathbf{a}$. The optimal form of $q_{n}$ is then given by:
$q_{\mathrm{opt}}(\mathbf{b})\propto p(\mathbf{b}|\theta_{n})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)$ (23)
@@
$q_{\mathrm{opt}}(\mathbf{b})=\frac{p(\mathbf{b}|\theta_{n})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)}{\int\mathrm{d}\mathbf{b}p(\mathbf{b}|\theta_{n})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)}$ (25)
@@ E.1 Derivation of $\mathcal{L}^{*}$
where $\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}$ is the Nyström approximation of $\mathbf{K}_{\mathbf{\hat{f}\hat{f}}}$. The trace will be small when $\mathbf{b}=\{f(\mathbf{x}_{n},\mathbf{a})\}$ and can be simplified as follows:
@@
Appendix F Online Upper Bound Implementation